OpenStudy (anonymous):

Write an equation of the line that passes through a pair of points: (-5,-2) (4,5)

OpenStudy (jennithemeani):

Start by finding the slope \[y^2-y^1/x^2-x^1\]

OpenStudy (jennithemeani):

y^2 is 5 and Y^1 is -2 X^2 is 4 and X^1 is -5 so \[5+2\div 4+5\]

OpenStudy (jennithemeani):

Add them together to get 7/9 This is your slope Now that you have your slope you have to find B So use y=mx+b and fill it in Take y and x from your equation (Make sure you pick the same y and x y^1 goes withX^1) 5=7/9(4)+b 5=28/9+b (mutiply 28/9 on both sides) 28/9 (5)= B B=140/9 So you then make it y=7/9(x)+140/9